OpenStudy (anonymous):

how to verify: sin 2x= 2 sinxcosx

OpenStudy (unklerhaukus):

\[\sin(2x)=\sin(x+x)=\sin(x)\cos(x)+\cos(x)\sin(x)=2\sin(x)\cos(x)\]

OpenStudy (anonymous):

yes!!